THE SPECTROSCOPY AND STRUCTURE OF THIOPHOSGENE (Cl_{2}CS) IN ITS SINGLET A_{2} STATE: A HIGH RESOLUTION STUDY OF THE ROTATIONAL STRUCTURE OF THE n → π^{∗} TRANSITION

T. FUJIWARA and E. C. LIM , Department of Chemistry and Centre for Laser and Optical Spectroscopy, The University of Akron, Akron, OH 44325-3601 ; \underline{R. H. JUDGE} and J. KODET, Department of Chemistry, University of Wisconsin-Parkside, Kenosha, WI 53141-2000 ; D. C. MOULE, Department of Chemistry, Brock University, St. Catharines, ON L2S 3A1.

BG test - The results of our recent line-by-line rotational analysis of the ring dye laser LIF jet spectrum of the ^{1}A_{2} ←^{1}A_{1} (n → π^{∗}) rovibronic transition of thiophosgene (Cl_{2}CS) will be presented. A total of four bands have been analyzed. An interesting feature of the spectrum is the axis-switching that occurs in the excited state of the ^{35}Cl_{2}CS and ^{35}Cl^{37}ClCS isotopomers. The excited state geometry has been determined from a fit to the least-squares determined rotational constants. The molecule is approximately 24^{∘} out-of-plane which is lower than that determined from earlier Franck-Condon type analyses of the vibrational structure. We attribut the discrepancy to the one-dimensional nature of the model calculations used in these previous studies.